OpenStudy (anonymous):

c(d)= 8 + 2d If Arelius wants to spend five days at the Senate forum, how many denari will he need for the trip? A. 18 B. 20 C. 22 D. 26

OpenStudy (anonymous):

Yay, Latin math :). Plug in 5 to your function: c(5) = 8+2*5. 8+2*5 = 18, so A.

OpenStudy (anonymous):

lets see if your right

OpenStudy (anonymous):

ummm is said i made an error for some reason it said it was wrong

OpenStudy (anonymous):

What is the total number of days for this trip? Remember, Arelius will spend two days traveling to Rome and two days traveling back from Rome.

OpenStudy (anonymous):

here is the whole question

OpenStudy (anonymous):

Arelius lives in a small town outside of Rome. He wants to see the Senate forum in Rome. The journey from his home to Rome is two days long and will cost 4 denari. Arelius calculates that with 2 denari per day, he will be able to pay for food and lodging. The function below can be used to determine c(d), the cost of the trip to Rome, where d represents the total number of days for the trip (including the days traveling to and from Rome). c(d)= 8 + 2d If Arelius wants to spend five days at the Senate forum, how many denari will he need for the trip?

OpenStudy (anonymous):

...well then, you would have to plug in 9 instead of 5.

OpenStudy (anonymous):

5 days to stay, 2 days to travel there, and 2 days to travel back.

OpenStudy (anonymous):

So c(9) = 8+2*9. 8+2*9 = 26, so D

OpenStudy (anonymous):

....scrossd was right based on the information that he was given, it's just that you didn't give him all the information, like the travel time back and forth....

OpenStudy (anonymous):

thank u for the answer i have many more

OpenStudy (anonymous):

....and how much denari would have to be spent each day. ._.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Actually, the answer should be 4 denari traveling to Rome, 2x5=10 denari for each day spent in the Senate, and another 4 denari traveling back from Rome. That should be.... 18 denari? Huh. Wait a second. ...shouldn't.... shouldn't 18 be right....? Huh. This is a strange question. The equation they gave doesn't really match with logic..... huh.

OpenStudy (anonymous):

That'w weird. They included the number of denari for the days traveling back and forth, but they still calculate it with the... 2 denari.... ohh I see. He still needs 2 denari for the days that he travels. I see.

OpenStudy (anonymous):

well when i checked the answer you were right 26 D

OpenStudy (anonymous):

Yeah, I forget that people need food and lodging even when they travel. :P So yeah, you would have to plug in 9. Sorry for the confusion....

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

So, you said you had other questions?